A motorboat maintained a constant speed of 14 miles per hour relative to the water in going 33 miles upstream and then returning. the total time for the trip was 7 hours. Use this information to find the speed of the current.
Answer by richwmiller(17219)

a+b=7,
(14+c)*a=33,
(14-c)*b=33
a = 3/2, b = 11/2, c = 8